### v1

# Nonreciprocal quantum phase transition in a spinning microwave magnonic system

Abstract
We propose how to achieve nonreciprocal quantum phase transition in a spinning microwave magnonic system composed of a spinning microwave resonator coupled with an yttrium iron garnet sphere with magnon Kerr effect. Sagnac-Fizeau shift caused by the spinning of the resonator brings about a significant modification in the critical driving strengths for second- and one-order quantum phase transitions, which means that the highly controllable quantum phase can be realized by the spinning speed of the resonator. More importantly, based on the difference in the detunings of the counterclockwise and clockwise modes induced by spinning direction of the resonator, the phase transition in this system is nonreciprocal, that is, the quantum phase transition occurs when the system is driven in one direction but not the other. Our work offers an alternative path to engineer and design nonreciprocal magnonic devices.
I Introduction
Quantum phase transition (QPT) has attracted considerable attention for its crucial role in quantum physics and potential applications in modern quantum technologies 1 ; 1a ; 2 ; 2a ; 2b ; 3 ; 4 ; 5 ; 6 ; 7 ; 9 ; 10 ; 11 ; 12 ; 13 ; 14 . Different from the classical phase transition happening in a finite temperature, this type of phase transition is mainly originated from the quantum fluctuation at the critical parameter regime and can take place between two stable phases accompanied by spontaneous symmetry breaking as the temperature tends to the absolute zero of temperature. Usually, a QPT occurs in the thermodynamic limit involving a infinite number of system components. Although recent advances about QPTs have indicated that the thermodynamic limit may be no longer a requirement 15 ; 16 ; 17 , the QPT observed in infinite-component system is still one of the most fascinating topics. Recently, it is interesting that that a squeezed light driving high-quality cavity can give rise to QPT in a dissipative qubit collective system without the ultra-strongly nonlinear coupling between the atoms and the field 18 . Additionally, the QPTs have also been studied in dissipative systems with other types of nonlinearity, such as the Kerr nonlinearity 19 ; 20 . Among them, it was confirmed that a QPT can emerge in a Kerr-nonlinear oscillator subject to two-photon driving and single-photon loss. Moreover, based on the magnon Kerr effect owing to the magnetocrystalline anisotropy in the ferromagnetic yttrium-iron-garnet sphere, the studies on QPT have made progress in hybrid cavity magnonic systems in recent years. In Ref. 21 , the authors studied the parity-symmetry-breaking QPT in a single-cavity magnonic system driven by a parametric field. This system can show a rich phase diagram, including the parity-symmetric phase (PSP), parity-symmetry-broken phase (PSBP), and bistable phase (BP). Subsequently, a controllable QPT was also proposed in a double-cavity magnonic system 22 . In particular, they demonstrated that the phase transition in one cavity can be precisely and efficiently controlled by adjusting the parameters of the other cavity. Nonetheless, whether the controllability of QPT can be realized only in a single-cavity magnonic system is still a question worth exploring.
Nonreciprocal physics, referring to the phenomenon that a system displays different responses in opposite directions, has witnessed rapid advances in recent years. Up to now, the study of nonreciprocity has been extended to many scientific branches, such as optics 23 ; 24 , acoustics 25 ; 26 ; 27 , thermodynamics 28 . As an important application, nonreciprocal devices allow signals to propagate in one direction and prevents signals from propagating in the opposite direction, and therefore plays a key role in backaction-immune communications. Traditional nonreciprocal devices break Lorentz reciprocity to achieve nonreciprocity mainly by employing the Faraday effect in magneto-optical crystal materials 29 ; 30 , but such devices needs the magnetic materials that are bulky, costly, and unsuitable for on-chip integration due to the high susceptibility to external magnetic field interference. To circumvent these obstacles, a variety of nonreciprocity schemes have been proposed based on nonlinear optics 31 ; 32 ; 32a ; 32b , optomechanics 33 ; 34 ; 35 , non-Hermitian optics 36 ; 37 ; 38 ; 39 . Especially, a recent experiment confirmed that an optical diode with 99.6% isolation has been realized by using a spinning resonator 40 , in which the nonreciprocity is caused by the Fizeau shift of circulating lights. Subsequently, there is increasing interests in the connection of the nonreciprocity with other quantum effects, such as nonreciprocal quantum blockades 41 ; 42 ; 42a ; 43 ; 44 ; 45 ; 46 ; 47 ; Wang22 ; 48 , quantum entanglement 49 ; 50 ; 51 ; 52 , mechanical squeezing 53 ; 54 ; 55 , phonon and magnon laser 56 ; 57 ; 58 , and sideband responses 59 ; 60 . Very recently, one particular concern is that nonreciprocal superradiant phase transitions have been reported in a cavity QED system 61 .
Here we propose how to achieve an nonreciprocal QPT in a single spinning microwave resonator coupled to an yttrium iron garnet (YIG) sphere with magnon Kerr nonlinearity. We show that, compared with the case of double-cavity, the QPT can be efficiently controlled only by adjusting the angular speed of the resonator without introducing an auxiliary resonator. Physically, the value of the order parameter for characterizing QPT is strongly dependent on the detuning between the cavity field and the driving laser. Meanwhile, the Sagnac effect caused by the spinning of the resonator can result in different influences on the detuning of the counterclockwise (CCW) and clockwise (CW) fields compared to the driving laser. Therefore, nonreciprocal QPT occurs by reason of the significant difference in the mean magnon number for driving the resonator from the opposite directions. Specifically, our system can display a rich phase diagram containing PSP, PSBP, and BP. Furthermore, we show that the system can experience a second-order phase transition from a PSP to a PSBP or a first-order phase transition from a PSP to a BP when increasing the driving strength beyond a critical value under different initial conditions. More importantly, in comparison to the stationary case, the tunable nonreciprocal QPT can be achieved by the Sagnac effect. Furthermore, we investigate the different behaviors of the mean magnon number and correlation fluctuation in the vicinity of the critical point and find that these behaviors can be used to identify different phase transitions. In addition, an isolation parameter is introduced to quantitatively describe the nonreciprocal degree. The results indicate again that the QPT can be readily switched on and off by adjusting the driving direction or the spinning direction of the resonator.
The remainder of this paper is organized as follows. In Sec. II, we introduce the theoretical model and give the dynamics and the steady-state solutions of the system through Heisenberg-Langevin approach. In Sec. III, the impact of the Sagnac-Fizeau shift and the rotating direction of the resonator on the nonreciprocal behaviors of the QPT are studied in detail. Finally, we summarize our results in Sec. IV.
II The model
Figure 1: (a) Schematic of the system. A YIG sphere with magnon Kerr nonlinearity is placed inside a spinning microwave resonator driven by a squeezed light generated by a flux-driven JPA and simultaneously biased by a uniform magnetic field along the z direction. (b) Frequencies of the microwave resonator. When the microwave resonator rotates in the fixed counterclockwise (CCW) direction, the resonance frequencies of the CCW mode and clockwise (CW) mode will experience different Fizeau shifts, where the frequencies $\omega_{a}+\left|\Delta_{F}\right|$ and $\omega_{a}-\left|\Delta_{F}\right|$ denote for the CCW and CW modes, respectively.
As schematically shown in Fig. 1, we consider a cavity-magnon system which is composed of a spinning microwave resonator coupled to a YIG sphere via the magnetic dipole interaction. While the resonator is driven by a weak squeezed vacuum field generated by a flux-driven Josephson parametric amplifier (JPA) and the magnetocrystalline anisotropy results in the Kerr nonlinear interaction among magnons in the YIG sphere. The Hamiltonian of the total system reads ($\hbar=1$)
$H$ $=$ $\left(\omega_{a}+\Delta_{F}\right)a^{\dagger}a+\omega_{m}m^{\dagger}m+\frac{K}{2}m^{\dagger}m^{\dagger}mm$ (1)
$+J\left(a^{\dagger}m+am^{\dagger}\right)+\frac{G}{2}\left(a^{\dagger 2}e^{-2i\omega_{d}t}+a^{2}e^{2i\omega_{d}t}\right),$
where $a$ $\left(a^{\dagger}\right)$ and $m$ $\left(m^{\dagger}\right)$ are, respectively, the annihilation (creation) operators of the cavity and magnon modes with the corresponding resonance frequencies $\omega_{a}$ and $\omega_{m}$. The microwave resonator is parametrically driven with driving amplitude $G$ and pump frequency $\omega_{d}$. $K$ denotes the nonlinear coefficient of the magnon Kerr effect and $J$ is the strength of the cavity-magnon interaction. For a resonator spinning at an angular velocity $\Omega$, the frequencies of the CCW and CW modes experience Sagnac-Fizeau shift, i.e., $\omega_{a}\rightarrow\omega_{a}+\Delta_{F}$, with
$\Delta_{F}=\pm\Omega\frac{nr\omega_{a}}{c}\left(1-\frac{1}{n^{2}}-\frac{\lambda}{n}\frac{dn}{d\lambda}\right),$ (2)
in which $n$ is the refractive index, $r$ is the radius of the resonator, and $\lambda$ $(c)$ is the wavelength (speed) of the light in vacuum. The dispersion term $dn/d\lambda$, characterizing the relativistic origin of the Sagnac effect, is relatively small ($\sim$1%) and thus can be ignored. In the rotating frame with respect to the driven frequency $\omega_{d}$, the Hamiltonian of the system becomes
$H$ $=$ $\tilde{\Delta}_{a}a^{\dagger}a+\Delta_{m}m^{\dagger}m+\frac{K}{2}m^{\dagger}m^{\dagger}mm$ (3)
$+J\left(a^{\dagger}m+am^{\dagger}\right)+\frac{G}{2}\left(a^{\dagger 2}+a^{2}\right),$
with $\tilde{\Delta}_{a}=\Delta_{a}+\Delta_{F},\Delta_{a}=\omega_{a}-\omega_{d}/2,\Delta_{m}=\omega_{m}-\omega_{d}/2$.
By considering dissipation and noise effects in the Hamiltonian (3), the quantum Langevin equations (QLEs) describing the dynamics of the hybrid system can be written as
$\dot{a}$ $=$ $-i\left(\tilde{\Delta}_{a}-i\kappa\right)a-iJm-iGa^{\dagger}+\sqrt{2\kappa}a_{in},$
$\dot{m}$ $=$ $-i\left(\Delta_{m}-i\gamma\right)m-iKm^{\dagger}mm-iJa+\sqrt{2\gamma}m_{in},$ (4)
where $\kappa$ $\left(\gamma\right)$ and $a_{in}$ $\left(m_{in}\right)$ are the damping rate and the zero-mean input noise operator for the cavity (magnon) mode, respectively. Furthermore, a linearization procedure can be used to simplify the physical model. We expand the operators $a$ and $m$ in Eq. (4) as a sum of their expectation values and small quantum fluctuations, $a=A+\delta a$ $\left(m=M+\delta m\right)$. By using the mean-field approximation, it follows from Eq. (4) that the dynamical equations for the expectation values are
$\dot{A}$ $=$ $-i\left(\tilde{\Delta}_{a}-i\kappa\right)A-iJM-iGA^{\ast},$
$\dot{M}$ $=$ $-i\left(\Delta_{m}-i\gamma\right)M-iK\left|M\right|^{2}M-iJA.$ (5)
In the steady-state case $\left(\dot{A}=\dot{M}=0\right)$, solving Eq. (5) gets three solutions for the mean magnon number $\left|M\right|^{2}$, i.e.,
$\left|M\right|_{0}^{2}=0,\text{ }\left|M\right|_{\pm}^{2}=\frac{-\tilde{\Delta}_{m}\pm\sqrt{\xi^{2}G^{2}-\tilde{\gamma}^{2}}}{K},$ (6)
with $\tilde{\Delta}_{m}=\Delta_{m}-\beta\tilde{\Delta}_{a}$, $\tilde{\gamma}=\gamma+\beta\kappa$, $\beta=J^{2}/\left(\tilde{\Delta}_{a}^{2}+\kappa^{2}-G^{2}\right)$. According to the first equation in Eq. (5), the steady-state photon occupation can be expressed as
$\left|A\right|_{0}^{2}=0,\text{ }\left|A\right|_{\pm}^{2}=\frac{\left(\tilde{\Delta}_{m}+\left|M\right|_{\pm}^{2}\right)^{2}+\gamma^{2}}{g^{2}}\left|M\right|_{\pm}^{2},$ (7)
hence we need only to consider the mean magnon number $\left|M\right|^{2}$ as an order parameter to characterize the QPT of the system. In order to derive critical conditions of the phase transition, we first concentrate our attention on $\left|M\right|_{+}^{2}\geq 0$. In the case of $-\tilde{\Delta}_{m}\geq 0$ (i.e., $\Delta_{m}/\tilde{\Delta}_{a}\leq\beta$), we take $\xi^{2}G^{2}-\tilde{\gamma}^{2}=0$ in Eq. (6), which gives rise to $G=G_{c1}$ with
$G_{c1}=\frac{-g^{2}+\sqrt{4\gamma^{2}\tilde{\Delta}_{a}^{2}+\left(g^{2}+2\gamma\kappa\right)^{2}}}{2\gamma}.$ (8)
Thus the critical driving strength $G_{c1}$ is independent of the detuning $\Delta_{m}$. In the other case, when $-\tilde{\Delta}_{m}\leq 0$ and $\sqrt{\xi^{2}G^{2}-\tilde{\gamma}^{2}}\geq 0$, the critical driving strength is
$G_{c2}=\sqrt{\frac{\left(g^{2}-\tilde{\Delta}_{a}\Delta_{m}\right)^{2}+\gamma^{2}\tilde{\Delta}_{a}^{2}+2g^{2}\gamma\kappa+\gamma^{2}\kappa^{2}+\Delta_{m}^{2}\kappa^{2}}{\gamma^{2}+\Delta_{m}^{2}}}.$ (9)
It clearly see from Eqs. (8) and (9) that both critical driving strengths are subject to the Fizeau shift $\Delta_{F}$. Secondly, we can use the nontrivial solution $\left|M\right|_{-}^{2}$ to obtain the third phase transition point even if it is proved to be unstable in the following. By only considering the case $-\tilde{\Delta}_{m}\geq 0$ and $\sqrt{\xi^{2}G^{2}-\tilde{\gamma}^{2}}\geq 0$, we obtain the critical driving strength $G=G_{c2}$, which has the same form as Eq. (9).
On the other hand, the linearized QLEs for quantum fluctuations can be written as
$\delta\dot{a}$ $=$ $-i\left(\tilde{\Delta}_{a}-i\kappa\right)\delta a-iJ\delta m-iG\delta a^{\dagger}+\sqrt{2\kappa}\delta a_{in},$
$\delta\dot{m}$ $=$ $-i\left(\tilde{\Delta}_{m}-i\gamma\right)\delta m-iJ\delta a-iKM^{2}\delta m^{\dagger}+\sqrt{2\gamma}\delta m_{in},$
with $\tilde{\Delta}_{m}=\Delta_{m}+2K\left|M\right|^{2}$. Here we have neglected the high-order terms of the fluctuations and $\delta o_{in}$ is the fluctuation on top of $o_{in}$ $\left(o=a,m\right)$. By defining the quadrature operators $\delta Q=\left(\delta a^{\dagger}+\delta a\right)/\sqrt{2}$, $\delta P=i\left(\delta a^{\dagger}-\delta a\right)/\sqrt{2}$, $\delta X=\left(\delta m^{\dagger}+\delta m\right)/\sqrt{2}$, $\delta Y=i\left(\delta m^{\dagger}-\delta m\right)/\sqrt{2}$, Eq. (LABEL:7) can be rewritten in a compact matrix form
$\delta\mathbf{\dot{O}}=\mathbf{U}\cdot\delta\mathbf{O}+\delta\mathbf{O}_{in},$ (11)
with the vector of quadrature components $\delta\mathbf{O=}\left(\delta Q,\delta P,\delta X,\delta Y\right)^{\text{T}}$ and the vector of noise quadratures $\delta\mathbf{O}_{in}=\left(\sqrt{2\kappa}\delta Q_{in},\sqrt{2\kappa}\delta P_{in},\sqrt{2\gamma}\delta X_{in},\sqrt{2\gamma}\delta Y_{in}\right)^{\text{T}}$. The superscript ‘T’ denotes the transpose. The drift matrix $\mathbf{U}$ takes the form
$\mathbf{U=}\left(\begin{array}[]{cccc}-\kappa&\Pi_{+}&0&J\\ \Pi_{-}&-\kappa&-J&0\\ 0&J&\Theta_{+}&\Xi_{+}\\ -J&0&\Xi_{-}&\Theta_{-}\end{array}\right),$ (12)
with $\Pi_{\pm}=\pm\tilde{\Delta}_{a}-G$, $\Theta_{\pm}=-\gamma\pm K$Im$\left[M^{2}\right]$ and $\Xi_{\pm}=\pm\tilde{\Delta}_{m}-K$Re$\left[M^{2}\right]$. For a given solution of $M$ in Eq. (5), only if all eigenvalues of the matrix $\mathbf{U}$ have negative real parts, the solution is said to be stable, otherwise this solution is unstable. After numerically carrying out this stability analysis [see Fig. 2(a)-2(c) and related discussions], we find that the nontrivial solution $\left|M\right|_{-}^{2}$ is always unstable in the whole parameter space, while the solutions $\left|M\right|_{-}^{2}$ and $\left|M\right|_{+}^{2}$ are stable in some parameter space, which suggests that we only need to apply the solutions $\left|M\right|_{0}^{2}$ and $\left|M\right|_{+}^{2}$ to investigate the steady-state quantum phase transitions. In addition, to research the variances of magnon mode quadratures, we define the time-dependent covariance matrix $V\left(t\right)$ with $V_{ij}\left(t\right)=\left\langle f_{i}\left(t\right)f_{j}\left(t^{\prime}\right)+f_{j}\left(t^{\prime}\right)f_{i}\left(t\right)\right\rangle/2$ $\left(i,j=1,2,3,4\right)$. From Eq. (11), one can easily find the solutions of $V\left(\infty\right)$ by solving the so-called Lyapunov equation $\mathbf{UV}+\mathbf{VU}^{T}=-\mathbf{D}$ with $\mathbf{D}$ being the diffusion matrix, defined as $\mathbf{D}_{ij}\left(t-t^{\prime}\right)=\left\langle\delta\mathbf{O}_{in,i}\left(t\right)\delta\mathbf{O}_{in,j}\left(t^{\prime}\right)+\delta\mathbf{O}_{in,j}\left(t^{\prime}\right)\delta\mathbf{O}_{in,i}\left(t\right)\right\rangle/2$. The diagonal elements of the matrix $\mathbf{V}$ associate with the variance of quadratures. In our model, the variance of magnon mode quadrature is given by $\left\langle\delta m^{\dagger}\delta m\right\rangle=\left[\left(\mathbf{V}_{33}+\mathbf{V}_{44}\right)-1\right]/2$.
III Nonreciprocal quantum phase transition
In this section, we explore how to realize an nonreciprocal QPT via the spinning of the resonator. We begin by employing the standard stability analysis to show the phase diagrams against the reduced driving strength $G/\kappa$ and the ratio $\Delta_{m}/\tilde{\Delta}_{c}$. As illustrated in Figs. 2(a)-2(c), each steady-state phase diagram exists three different regions corresponding to different phases, i.e., PSP, PSBP, and BP. The vertical black solid line ($G=G_{c1}$) is the boundary between PSP and BP in the case of $\Delta_{m}/\tilde{\Delta}_{a}\leq\beta|_{G=Gc1}$, and then the blue and red solid curves ($G=G_{c2}$) are the boundaries between PSP and PSBP for $\Delta_{m}/\tilde{\Delta}_{a}\geq\beta|_{G=Gc1}$, BP and PSBP for $\Delta_{m}/\tilde{\Delta}_{a}\leq\beta|_{G=Gc1}$, respectively. These manifest that the boundaries between different phases are determined by the critical driving strengths $G_{c1}$ and $G_{c2}$. In particular, the intersection point of three boundaries emerges when $G_{c1}=G_{c2}$ at $\Delta_{m}/\tilde{\Delta}_{a}=\beta|_{G=Gc1}$. More interestingly, we find that different spinning directions (the left or right) give rise to QPT at different critical driving strength. Compared with the stationary case (i.e., no spinning with $\Delta_{F}=0$), the areas of three phases have a remarkable change for $\Delta_{F}>0$ and $\Delta_{F}<0$. Therefore, this QPT are highly controllable and can be tuned by the driving direction (or the spinning direction) and the spinning speed of the resonator. As a result, the nonreciprocal QPT can be achieved in this system.
Figure 2: Steady-state phase diagram of the system varying with the reduced driving strength $G/\kappa$ and the detuning ratio $\Delta_{m}/\tilde{\Delta}_{c}$, in which PSP, PSBP, and BP represent the areas for the parity-symmetric phase, parity-symmetry-broken phase and bistable phase, respectively. The Fizeau shifts are $\Delta_{F}/\kappa=-0.3$ in (a), $\Delta_{F}/\kappa=0$ in (b), and $\Delta_{F}/\kappa=0.3$ in (c). The other parameters are $\Delta_{c}/\kappa=3$, $J/\kappa=2.5$, and $\gamma/\kappa=1.$
In Figs. 3(a)-2(c), we further display the dynamical behaviors of the scaled mean magnon number $\left|M\right|^{2}/\left(\gamma/K\right)$ to reveal the macroscopic magnon excitations for different phases. Here the parameter values for each phase have marked with different-colored dots in Fig. 2(b). When the system is in BP [see Fig. 3(a)], we find that both the solutions $\left|M\right|_{0}^{2}$ and $\left|M\right|_{+}^{2}$ are stable and the steady-state magnon occupation can be dominated by the initial condition of the system. This situation indicates that the parity symmetry in BP can be conserved or broken with appropriate initial conditions. It is noticed that PSP and BP will cannot be distinguished by only tracking the dynamics of the system with an initial state near zero since the solution $\left|M\right|_{0}^{2}$ is stable for both phases. In this case, the initial state should be set far away from zero to distinguish the two phases and observe the phase transition. From Fig. 3(b), we see that only the nontrivial solution $\left|M\right|_{+}^{2}$ becomes stable in PSBP. The system has the macroscopic magnon excitation in this phase and thus is parity-symmetry broken. As described in Fig. 3(c), there is no macroscopic magnon excitation in PSP, that is to say, only the solution $\left|M\right|_{0}^{2}$ is stable and then the system is parity symmetric. Besides, an interesting feature is that the nontrivial solution $\left|M\right|_{-}^{2}$ is always unstable in the whole parameter range since the four eigenvalues of the matrix $\mathbf{U}$ in Eq. (12) have at least one nonnegative real part.
Figure 3: Dynamics of the normalized magnon number for three phases at the points marked by different-colored dots in Fig. 2(b), with (a) $G/\kappa=1.7$ and $\Delta_{m}/\Delta_{c}=1$, (b) $G/\kappa=2.3$ and $\Delta_{m}/\Delta_{c}=1.2$, (c) $G/\kappa=2.3$ and $\Delta_{m}/\Delta_{c}=0.4$. In (a)-(c), $\left\langle a\right\rangle_{t=0}/\sqrt{\gamma/K}=2+1.5i$ and $\left\langle b\right\rangle_{t=0}/\sqrt{\gamma/K}=2-1.5i$ for the solid curves, while $\left\langle a\right\rangle_{t=0}/\sqrt{\gamma/K}=0.05+0.05i$ and $\left\langle b\right\rangle_{t=0}/\sqrt{\gamma/K}=0.05-0.05i$ for the dashed curves. The other parameters are $J/\kappa=2.5$ and $\gamma/\kappa=1.$
[…]
Figure 4: Normalized steady-state magnon number $\left|M\right|^{2}/\left(\gamma/K\right)$ is plotted as the scaled driving strength $G/\kappa$ for three different Fizeau shifts $\Delta_{F}/\kappa=-0.3$, 00, and $0.3$. The solid curves correspond to the analytical results in Eq. (6) and the circles correspond to the numerical results obtained by Eq. (LABEL:7). (a) $\Delta_{m}/\kappa=4$ and initial conditions $\left\langle a\right\rangle_{t=0}/\sqrt{\gamma/K}=\left\langle b\right\rangle_{t=0}/\sqrt{\gamma/K}=0.2+0.2i$, (b) $\Delta_{m}/\kappa=2.2$ and initial conditions $\left\langle a\right\rangle_{t=0}/\sqrt{\gamma/K}=\left\langle b\right\rangle_{t=0}/\sqrt{\gamma/K}=10+10i$. The other parameters are $\Delta_{c}/\kappa=3$, $J/\kappa=2.5$, and $\gamma/\kappa=1.$
Next, we reveal more characteristics of the nonreciprocal QPT through numerically observing the behavior of the correlated fluctuation of the magnon operators $\delta m^{\dagger}$ and $\delta m$. As shown in Fig. 5, we plot the expectation value $\left\langle\delta m^{\dagger}\delta m\right\rangle$ of correlated fluctuation as a function of the scaled coupling strength $G/\kappa$ for different values of $\Delta_{F}$. There is a common characteristic in both Figures 5(a) and (5b), that is, the mean correlated fluctuation $\left\langle\delta m^{\dagger}\delta m\right\rangle$ is close to zero when the driving strength $G$ is away from the critical value, but Figure 5(a) (Figure 5(b)) is of a asymptotic (catastrophic) divergent behavior around the critical point, which is connected with the second-order QPT from PSP to PSBP (the first-order QPT from PSP to BP). From Fig. 5, we notice especially that spinning the resonator decreases the critical driving strength for $\Delta_{F}<0$ or increases it for $\Delta_{F}>0$, compared with the stationary-resonator case ($\Delta_{F}=0$). Besides, associating Fig. 4 with Fig. 5 finds that the the critical positions where the QPT occurs in Fig. 5(a) (Fig. 5(b)) completely agree with those in Fig. 4(a) (Fig. 4(b)) as we expected. The results give further evidence that the nonreciprocal QPT can be realized in the model.
Figure 5: Mean correlated fluctuation $\left\langle\delta m^{\dagger}\delta m\right\rangle$ as a function of the scaled driving strength $G/\kappa$ for three different Fizeau shifts $\Delta_{F}/\kappa=-0.3,0,$and $0.3$. The other parameters are the same with Figure 4.
Finally, to quantitatively describe the effect of spinning on the nonreciprocal QPT, we introduce the isolation parameter
=$\mathfrak{R}=\left\{\begin{array}[]{cc}0,\text{ }&\left|M\right|^{2}\left(\Delta_{F}\right)=0\\ \left|\frac{\left|M\right|^{2}\left(\Delta_{F}<0\right)-\left|M\right|^{2}\left(\Delta_{F}>0\right)}{\left|M\right|^{2}\left(\Delta_{F}<0\right)+\left|M\right|^{2}\left(\Delta_{F}>0\right)}\right|.&The\text{ }other\end{array}\right.$ (13)
For a QPT without the spinning resonator (i.e., a conventional reciprocal QPT), the isolation parameter is $\mathfrak{R}=0$. A nonzero $\mathfrak{R}$ denotes the emergence of nonreciprocity in the phase transition. The higher the isolation parameter $\mathfrak{R}$ is, the stronger the nonreciprocity of the QPT is. Especially, $\mathfrak{R}=1$ corresponds to an ideal nonreciprocal QPT. The isolation parameter $\mathfrak{R}$ as the function of the Fizeau shift $\left|\Delta_{F}\right|/\kappa$ is shown Figs. 6(a) and 6(b) for the second-order and first-order phase transitions, respectively. As expected, the isolation parameter is $\mathfrak{R}=0$ for the QPT with a stationary resonator (i.e., $\Delta_{F}=0$). While it is clear that, not only the nonreciprocal QPT occurs in a remarkably broad parameter range, but also the region of the ideal nonreciprocal QPT enlarges with the increasing of the Fizeau shift $\left|\Delta_{F}\right|$. Further, the ideal nonreciprocal QPT exists obviously two boundaries, one boundary condition (the left black curve) satisfies $G=G_{c2}\left(\Delta_{F}<0\right)$ and the other one (the right black curve) is $G=G_{c2}\left(\Delta_{F}>0\right)$ in Fig. 5(a). In the same way, there exists a similar characteristic in Fig. 5(b). An interesting feature in Fig. 6 is that in the vicinity of the left black curses, the isolation parameter $\mathfrak{R}$ experiences a sudden change from zero to one. That is because both mean magnon numbers $\left|M\right|^{2}\left(\Delta_{F}<0\right)$ and $\left|M\right|^{2}\left(\Delta_{F}>0\right)$ are zero when the driving strength is less than the critical strength, However, when the driving strength is more than the the critical strength, $\left|M\right|^{2}\left(\Delta_{F}<0\right)$ is nonzero and $\left|M\right|^{2}\left(\Delta_{F}>0\right)$ is still zero. In this situation, we see that the isolation parameter $\mathfrak{R}$ undergoes a gradually change from one to zero. It can be well understood that $\left|M\right|^{2}\left(\Delta_{F}<0\right)$ is nonzero and $\left|M\right|^{2}\left(\Delta_{F}>0\right)$ is zero when the driving strength is less than the critical strength, on the other hand, when the driving strength is more than the critical strength, both $\left|M\right|^{2}\left(\Delta_{F}<0\right)$ and $\left|M\right|^{2}\left(\Delta_{F}>0\right)$ are not only nonzero, but also gradually increasing with the driving strength increasing.
Figure 6: Isolation parameter $\mathfrak{R}$ versus the scaled Fizeau shift $\left|\Delta_{F}\right|/\kappa$ and normalized driving strength $G/\kappa$. The left and right black curves correspond to the boundary conditions in Eqs. (8) and (9), respectively. The parameters are the same as those in Figure 2.
IV Conclusions
In summary, we have theoretically studied the role of rotation in engineering an nonreciprocal quantum phase transition in the cavity magnonic system consisting of a YIG sphere coupled to a spinning microwave resonator. By rotating the microwave resonator, we show that the introduced Sagnac effect can significantly modify the critical driving strengths for both second- and first-order quantum phase transitions and lead to a largely adjustable range of the critical driving strength. We further find that the critical driving strength of the phase transition relies on both the driving direction (or the rotation direction) and the spinning speed. In consequence, a highly tunable nonreciprocal quantum phase transition can be achieved based on the Sagnac effect. Our proposal provides an alternative route to achieve nonreciprocal quantum phase in microwave magnonical system and may find promising applications in designing nonreciprocal magnonic devices.
Acknowledgments
This work is supported by the National Science Foundation for Distinguished Young Scholars of the Higher Education Institutions of Anhui Province (Grant No. 2022AH020097). G. Q. Zhang is supported by the National Natural Science Foundation of China (Grant No. 12205069).
Note added. -In preparing our manuscript, we became aware of a similar work on nonreciprocal superradiant phase transitions published on Physical Review Letters 61 .