### v2

# Nonreciprocal quantum phase transition in cavity magnonics

Abstract
We investigate the nonreciprocal quantum phase transition in a cavity magnonic system driven by a parametric field, where an yttrium iron garnet (YIG) sphere is placed in a spinning microwave resonator. The system exhibits a rich phase diagram due to both magnon Kerr nonlinearity in YIG and parametric drive on the resonator. Especially, Sagnac-Fizeau shift caused by the spinning of the resonator brings about a significant modification in the critical driving strengths for second- and first-order quantum phase transitions, which means that the highly controllable quantum phase can be realized by the spinning speed of the resonator. More importantly, based on the difference in the detunings of the counterclockwise and clockwise modes induced by spinning direction of the resonator, we show that the phase transition in this system is nonreciprocal, that is, the quantum phase transition occurs when the cavity is driven in one direction but not the other. Our work offers an alternative path to engineer and design nonreciprocal magnonic devices.
I Introduction
In recent years, the cavity magnonic system has become an outstanding platform for exploring light-matter interactions Lachance-Quirion2019 ; Rameshti2022 ; PengYan-2021 . Since ferrimagnetic systems, e.g., yttrium iron garnet (YIG), own high spin density and low damping rate, the strong (even ultra-strong) coupling between magnons in ferrimagnetic systems and photons in microwave cavities can be achieved experimentally Huebl2013 ; Tabuchi2014 ; Zhang2014 ; Goryachev2014 ; Hu-PRL-15 ; Zhang-npjQI-15 . This strong coherent interaction allows one to study many fascinating phenomena, such as magnon dark modes 01 , magnon-photon entanglement 01a , magnon blockade 01b ; 01c ; 01d , non-Hermitian physics 02 ; 01e ; Zhang19 ; 01f ; 01g , quantum states of magnons 01h ; 01i ; 01j ; Zhang23 , cooperative polariton dynamics 01k , and magnon spintronics 01n ; 03 . In addition, it is worth mentioning that the magnon Kerr effect, originating from the magnetocrystalline anisotropy in the YIG, has been theoretically and experimentally demonstrated in cavity-magnon systems 03a ; 03b . The magnon Kerr effect gives rise to the development of some topics, including bistability and tristability of cavity-magnon polaritons 04 ; Nair20 ; Bi21 ; Shen21 ; Bi24 , high-order sideband generation 05 , coherent perfect absorption Zhang-Wang23 , quantum entanglement Yang21 ; Zhang-Agarwal19 , and long-range spin-spin coupling 08 .
Owing to its crucial role in quantum physics and potential applications in quantum technologies, the quantum phase transition (QPT) has attracted considerable attention1 ; 1a ; 2 ; 2a ; 2b ; 3 ; 4 ; 5 ; 6 ; 7 ; 9 ; 10 ; 11 ; 12 ; 13 ; 14 . Different from the classical phase transition happening at a finite temperature, the QPT mainly originates from the quantum fluctuation. In the critical parameter regime, the QPT occurs between two stable phases accompanied by spontaneous symmetry breaking as the temperature tends to the absolute zero of temperature. Recently, QPT has not only been investigated in the Dicke model (or the Tavis-Cummings model) without the ultrastrongly coupling requirement when the cavity field is squeezed via parametric drive 18 , but also been studied in dissipative systems with other types of nonlinearity (e.g., Kerr nonlinearity) 19 ; 20 . With the assistance of magnon Kerr effect and parametric drive, the QPT has also been explored in cavity magnonics 21 ; 22 .
Nonreciprocal physics refers to the phenomenon that a system displays different responses in opposite directions. Up to now, the study of nonreciprocity has been extended to many scientific branches, such as optics 23 ; 24 , acoustics 25 ; 26 ; 27 , and thermodynamics 28 . Traditional nonreciprocal devices break Lorentz reciprocity to achieve nonreciprocity mainly by employing the Faraday effect in magneto-optical crystal materials 29 ; 30 . In such nonreciprocal devices, the used magnetic materials are bulky due to the requirement of high susceptibility to external magnetic field interference, which makes them unsuitable for on-chip integration. To circumvent these obstacles, a variety of nonreciprocity schemes have been proposed based on nonlinear optics 31 ; 32 ; 32a ; 32b , optomechanics 33 ; 34 ; 35 , non-Hermitian optics 36 ; 37 ; 38 ; 39 , etc. In particular, a recent experiment confirmed that an optical diode with 99.6% isolation has been realized by using a spinning resonator 40 , in which the nonreciprocity is caused by the Fizeau shift of circulating lights. Subsequently, there are increasing interests in the connection of the nonreciprocity with other quantum effects, such as quantum blockades 41 ; 42 ; 42a ; 43 ; 44 ; 45 ; 46 ; 47 ; 47a ; 48 , quantum entanglement 49 ; 50 ; 51 ; Chen23 ; 52 , mechanical squeezing 53 ; 54 ; 55 , phonon and magnon laser 56 ; 57 ; 58 , sideband responses 59 ; 60 , single-photon state conversion 60a , and superradiant phase transitions 61c ; 61d ; 61 . However, the nonreciprocal QPT has not yet been investigated in cavity magnonics.
Here we propose a scheme to realize a nonreciprocal QPT in a spinning microwave magnonic system. We first display the phase diagram of the hybrid system, which contains parity-symmetric phase (PSP), parity-symmetry-broken phase (PSBP), and bistable phase (BP). In particular, the Sagnac effect caused by the spinning of the resonator can result in different influences on the detuning of the counterclockwise (CCW) and clockwise (CW) fields compared to the driving field. This leads to the occurrence of the nonreciprocal QPT for driving the resonator from the opposite directions. By investigating the different behaviors of the mean magnon number in the vicinity of the critical point for different Fizeau shifts, we show that the QPT is not only either continuous (second-order) or discontinuous (first-order), but also nonreciprocal for different driving directions. Moreover, an isolation parameter is introduced to quantitatively describe the nonreciprocal QPT. At last the nonreciprocal fluctuation of magnon number is also numerically simulated for displaying the features of the nonreciprocal QPT from the point of view of quantum fluctuation. Nonreciprocal QPT may open up the prospect of engineering nonreciprocal devices for applications in, e.g., on-chip unidirectional quantum sensing 61a and quantum metrology 61b .
In other systems, nonreciprocal phase transition has been investigated 61c ; 61d ; 61 . Especially, Ref. 61c makes a significant contribution to the theory of nonreciprocal phase transition, which originates from asymmetric interactions of multiple species. Subsequently, nonreciprocal coupling is further used to realize nonreciprocal QPT in an open Dicke model 61d . Very recently, Zhu et al 61 proposed to study nonreciprocal QPT in an open dual-coupling Jaynes-Cummings model by utilizing Sagnac effect, where a two-level atom is nonlinearly coupled to modes of a spinning microcavity. As an advantage of this work, it is demonstrated that in the presence of cavity dissipation, the squeezed light can recover the QPT through breaking the requirement of ultrastrong atom-field coupling. On the contrary, although we propose to achieve the nonreciprocity of QPT based on Sagnac effect, our scheme is established in a spinning microwave magnonical system. Here the achievement of QPT relies on not only the squeezed field, but also the magnon Kerr nonlinearity in the YIG. Our scheme exploits a different mechanism to engineer nonreciprocal QPT.
The remainder of this paper is organized as follows. In Sec. II, we introduce the theoretical model. Using the Heisenberg-Langevin approach, we give the steady-state solutions of the system. In Sec. III, the nonreciprocal QPT, induced by the Sagnac-Fizeau shift, is studied in detail. Finally, we summarize our results in Sec. IV.
II The model
Figure 1: (a) Schematic of the system. A YIG sphere, biased by a uniform magnetic field $B_{z}$ along the $z$ direction, is placed inside a spinning microwave resonator driven by a squeezed light generated by a flux-driven JPA. (b) Frequencies of the microwave resonator. When the microwave resonator rotates in the fixed counterclockwise (CCW) direction, the resonance frequencies of the CCW mode and clockwise (CW) mode will experience different Fizeau shifts, where the frequencies $\omega_{a}+\left|\Delta_{F}\right|$ and $\omega_{a}-\left|\Delta_{F}\right|$ denote for the CCW and CW modes, respectively.
As schematically shown in Fig. 1, we consider a cavity-magnon system consisting of a spinning resonator and a YIG sphere, where photons in the resonator are coupled to magnons in YIG. In the hybrid system, the resonator is driven by a weak squeezed vacuum field generated by a flux-driven Josephson parametric amplifier (JPA) and the magnetocrystalline anisotropy of YIG results in the Kerr nonlinear interaction among magnons. The Hamiltonian of the total system reads ($\hbar=1$)
$H$ $=$ $\left(\omega_{a}+\Delta_{F}\right)a^{\dagger}a+\omega_{m}m^{\dagger}m+\frac{K}{2}m^{\dagger}m^{\dagger}mm$ (1)
$+J\left(a^{\dagger}m+am^{\dagger}\right)+\frac{G}{2}\left(a^{\dagger 2}e^{-2i\omega_{d}t}+a^{2}e^{2i\omega_{d}t}\right),$
where $a$ $\left(a^{\dagger}\right)$ and $m$ $\left(m^{\dagger}\right)$ are, respectively, the annihilation (creation) operators of the cavity and magnon modes with the corresponding resonance frequencies $\omega_{a}$ and $\omega_{m}$. The microwave resonator is parametrically driven with driving amplitude $G$ and pump frequency $\omega_{d}$. $K$ denotes the nonlinear coefficient of the magnon Kerr effect and $J$ is the strength of the cavity-magnon interaction. For a resonator spinning at an angular velocity $\Omega$, the frequencies of the CCW and CW modes experience Sagnac-Fizeau shift, i.e., $\omega_{a}\rightarrow\omega_{a}+\Delta_{F}$, with 40 ; 62
$\Delta_{F}=\pm\Omega\frac{nr\omega_{a}}{c}\left(1-\frac{1}{n^{2}}-\frac{\lambda}{n}\frac{dn}{d\lambda}\right),$ (2)
in which $n$ is the refractive index, $r$ is the radius of the resonator, and $\lambda$ $(c)$ is the wavelength (speed) of the light in vacuum. The dispersion term $dn/d\lambda$, characterizing the relativistic origin of the Sagnac effect, is relatively small ($\sim$1%) and thus can be ignored.
In the rotating frame with respect to the driven frequency $\omega_{d}$, the Hamiltonian of the system becomes
$\mathcal{H}$ $=$ $\left(\tilde{\Delta}_{a}-i\kappa\right)a^{\dagger}a+\left(\Delta_{m}-i\gamma\right)m^{\dagger}m+\frac{K}{2}m^{\dagger}m^{\dagger}mm$ (3)
$+J\left(a^{\dagger}m+am^{\dagger}\right)+\frac{G}{2}\left(a^{\dagger 2}+a^{2}\right),$
where the dissipations of the cavity-magnon system have been considered. Here $\kappa$ ($\gamma$) is the damping rate of the photon (magnon) mode, $\tilde{\Delta}_{a}=\Delta_{a}+\Delta_{F}$, $\Delta_{a}=\omega_{a}-\omega_{d}/2$, and $\Delta_{m}=\omega_{m}-\omega_{d}/2$. Note that in the absence of YIG sphere (i.e., $J=0$), the squeezing driving can cause the system unstable if $G>\sqrt{\tilde{\Delta}_{a}^{2}+\kappa^{2}}$ 63 . Hence we only consider the case of $G<\sqrt{\tilde{\Delta}_{a}^{2}+\kappa^{2}}$. Interestingly, the non-Hermitian Hamiltonian $\mathcal{H}$ in Eq. (3) commutes with the parity operator $\mathcal{P}=\exp\left[i\pi\left(a^{\dagger}a+m^{\dagger}m\right)\right]$ 2 , i.e., $\left[\mathcal{H},\mathcal{P}\right]=0$, which indicates that the Hamiltonian $\mathcal{H}$ has the parity symmetry. Normally, the total excitation number of the system at steady states can be any natural number. When it takes odd (even) numbers, the steady state of the system has the odd (even) parity. If the total excitation number of the system versus system parameters varies in the even subset $\{0,2,4,6,8,\cdots\}$ or the even subset $\{1,3,5,7,9,\cdots\}$, we say that the corresponding steady state of the system is parity-symmetric. Otherwise, the steady state of the system is parity-symmetry-breaking.
By considering dissipation and noise effects in the Hamiltonian (3) 64 , the quantum Langevin equations (QLEs) describing the dynamics of the hybrid system can be written as
$\dot{a}$ $=$ $-i\left(\tilde{\Delta}_{a}-i\kappa\right)a-iJm-iGa^{\dagger}+\sqrt{2\kappa}a_{\text{in}},$
$\dot{m}$ $=$ $-i\left(\Delta_{m}-i\gamma\right)m-iKm^{\dagger}mm-iJa+\sqrt{2\gamma}m_{\text{in}},$ (4)
where $a_{\text{in}}$ and $m_{\text{in}}$ are the zero-mean input noise operator for the microwave photon (magnon) mode, respectively. To linearize the Eq. (4), we can expand the operators $a$ and $m$ as a summation of their expectation values and quantum fluctuations, i.e., $a=A+\delta a$ and $m=M+\delta m$. Following from Eq. (4), the dynamical equations for the expectation values $A$ and $M$ are
$\dot{A}$ $=$ $-i\left(\tilde{\Delta}_{a}-i\kappa\right)A-iJM-iGA^{\ast},$
$\dot{M}$ $=$ $-i\left(\Delta_{m}-i\gamma\right)M-iK\left|M\right|^{2}M-iJA.$ (5)
In the steady-state case ($\dot{A}=\dot{M}=0$), solving Eq. (5) gets three solutions for the mean magnon number $\left|M\right|^{2}$, i.e.,
$\left|M\right|_{0}^{2}=0,\text{ }\left|M\right|_{\pm}^{2}=\frac{-\tilde{\Delta}_{m}\pm\sqrt{\beta^{2}G^{2}-\tilde{\gamma}^{2}}}{K},$ (6)
with $\tilde{\Delta}_{m}=\Delta_{m}-\beta\tilde{\Delta}_{a}$, $\tilde{\gamma}=\gamma+\beta\kappa$, and $\beta=J^{2}/$($\tilde{\Delta}_{a}^{2}+\kappa^{2}-G^{2}$). According to the first equation in Eq. (5), the steady-state photon occupation can be expressed as
$\left|A\right|_{0}^{2}=0,\text{ }\left|A\right|_{\pm}^{2}=\frac{\left(\tilde{\Delta}_{m}+\left|M\right|_{\pm}^{2}\right)^{2}+\gamma^{2}}{g^{2}}\left|M\right|_{\pm}^{2}.$ (7)
Thus, we need only to consider the mean magnon number $\left|M\right|^{2}$ as an order parameter to characterize the QPT of the system. In order to derive critical conditions of the phase transition, we first concentrate our attention on $\left|M\right|_{+}^{2}\geq 0$. In the case of $-\tilde{\Delta}_{m}\geq 0$ (i.e., $\Delta_{m}/\tilde{\Delta}_{a}\leq\beta$), we take $\beta^{2}G^{2}-\tilde{\gamma}^{2}=0$ in Eq. (6), which gives rise to $G=G_{\text{c1}}$ with
$G_{\text{c1}}=\frac{-g^{2}+\sqrt{4\gamma^{2}\tilde{\Delta}_{a}^{2}+\left(g^{2}+2\gamma\kappa\right)^{2}}}{2\gamma}.$ (8)
Thus the critical driving strength $G_{\text{c1}}$ is independent of the detuning $\Delta_{m}$. In the other case, when $-\tilde{\Delta}_{m}\leq 0$ and $\sqrt{\beta^{2}G^{2}-\tilde{\gamma}^{2}}\geq 0$, the critical driving strength is
$G_{\text{c2}}=\sqrt{\frac{\left(g^{2}-\tilde{\Delta}_{a}\Delta_{m}\right)^{2}+\gamma^{2}\tilde{\Delta}_{a}^{2}+2g^{2}\gamma\kappa+\gamma^{2}\kappa^{2}+\Delta_{m}^{2}\kappa^{2}}{\gamma^{2}+\Delta_{m}^{2}}}.$ (9)
It clearly see from Eqs. (8) and (9) that both critical driving strengths are subject to the Fizeau shift $\Delta_{F}$. Secondly, we can use the nontrivial solution $\left|M\right|_{-}^{2}$ to obtain the third phase transition point even if it is proved to be unstable in the following. By only considering the case $-\tilde{\Delta}_{m}\geq 0$ and $\sqrt{\beta^{2}G^{2}-\tilde{\gamma}^{2}}\geq 0$, we obtain the critical driving strength $G=G_{\text{c2}}$, which has the same form as Eq. (9).
On the other hand, the linearized QLEs for quantum fluctuations can be written as
$\delta\dot{a}$ $=$ $-i\left(\tilde{\Delta}_{a}-i\kappa\right)\delta a-iJ\delta m-iG\delta a^{\dagger}+\sqrt{2\kappa}a_{\text{in}},$
$\delta\dot{m}$ $=$ $-i\left(\tilde{\Delta}_{m}-i\gamma\right)\delta m-iJ\delta a-iKM^{2}\delta m^{\dagger}+\sqrt{2\gamma}m_{\text{in}},$
with $\tilde{\Delta}_{m}=\Delta_{m}+2K\left|M\right|^{2}$. Here we have neglected the high-order terms of the fluctuations. By defining the quadrature operators $\delta Q=\left(\delta a^{\dagger}+\delta a\right)/\sqrt{2}$, $\delta P=i\left(\delta a^{\dagger}-\delta a\right)/\sqrt{2}$, $\delta X=\left(\delta m^{\dagger}+\delta m\right)/\sqrt{2}$, and $\delta Y=i\left(\delta m^{\dagger}-\delta m\right)/\sqrt{2}$, Eq. (LABEL:7) can be rewritten in a compact matrix form
$\delta\mathbf{\dot{O}}=\mathbf{U}\cdot\delta\mathbf{O}+\mathbf{O}_{\text{in}},$ (11)
with the vector of quadrature components $\delta\mathbf{O=}\left(\delta Q,\delta P,\delta X,\delta Y\right)^{\text{T}}$ and the vector of noise quadratures $\mathbf{O}_{\text{in}}=\left(\sqrt{2\kappa}Q_{\text{in}},\sqrt{2\kappa}P_{\text{in}},\sqrt{2\gamma}X_{\text{in}},\sqrt{2\gamma}Y_{\text{in}}\right)^{\text{T}}$ with $Q_{\text{in}}=\left(a_{\text{in}}^{\dagger}+a_{\text{in}}\right)/\sqrt{2}$, $P_{\text{in}}=i\left(a_{\text{in}}^{\dagger}-a_{\text{in}}\right)/\sqrt{2}$, $X_{\text{in}}=\left(m_{\text{in}}^{\dagger}+m_{\text{in}}\right)/\sqrt{2}$, and $Y_{\text{in}}=i\left(m_{\text{in}}^{\dagger}-m_{\text{in}}\right)/\sqrt{2}$. The superscript ‘T’ denotes the transpose. The drift matrix $\mathbf{U}$ takes the form
$\mathbf{U=}\left(\begin{array}[]{cccc}-\kappa&\Pi_{+}&0&J\\ \Pi_{-}&-\kappa&-J&0\\ 0&J&\Theta_{+}&\Xi_{+}\\ -J&0&\Xi_{-}&\Theta_{-}\end{array}\right),$ (12)
with $\Pi_{\pm}=\pm\tilde{\Delta}_{a}-G$, $\Theta_{\pm}=-\gamma\pm K$Im$\left[M^{2}\right]$ and $\Xi_{\pm}=\pm\tilde{\Delta}_{m}-K$Re$\left[M^{2}\right]$. For a given solution of $M$ in Eq. (5), only if all eigenvalues of the matrix $\mathbf{U}$ have negative real parts, the solution is said to be stable, otherwise this solution is unstable 65 . After numerically carrying out this stability analysis [see Fig. 2 and related discussions], we find that the nontrivial solution $\left|M\right|_{-}^{2}$ is always unstable in the whole parameter space, while the solutions $\left|M\right|_{0}^{2}$ and $\left|M\right|_{+}^{2}$ are stable in some parameter space, which suggests that we only need to apply the solutions $\left|M\right|_{0}^{2}$ and $\left|M\right|_{+}^{2}$ to investigate the steady-state quantum phase transition. In addition, to research the variances of magnon mode quadratures, we define the time-dependent covariance matrix $\mathbf{V}\left(t\right)$ with $V_{\text{ij}}\left(t\right)=\left\langle f_{i}\left(t\right)f_{j}\left(t^{\prime}\right)+f_{j}\left(t^{\prime}\right)f_{i}\left(t\right)\right\rangle/2$ $\left(i,j=1,2,3,4\right)$. From Eq. (11), one can easily find the solutions of $V\left(\infty\right)$ by solving the so-called Lyapunov equation $\mathbf{UV}+\mathbf{VU}^{T}=-\mathbf{D}$ with $\mathbf{D}$ being the diffusion matrix, defined as $D_{\text{ij}}\left(t-t^{\prime}\right)=\left\langle\delta O_{\text{in},i}\left(t\right)\delta O_{\text{in},j}\left(t^{\prime}\right)+\delta O_{\text{in},j}\left(t^{\prime}\right)\delta O_{\text{in},i}\left(t\right)\right\rangle/2$. The diagonal elements of the matrix $\mathbf{V}$ associate with the variance of quadratures. In our model, the variance of magnon mode quadrature is given by $\langle\delta m^{\dagger}\delta m\rangle$=$\left[\left(V_{33}+V_{44}\right)-1\right]/2$ 18 .
III Nonreciprocal quantum phase transition
In this section, we explore how to realize a nonreciprocal QPT via the spinning of the resonator. We begin by employing the standard stability analysis to show the phase diagrams against the reduced driving strength $G/\kappa$ and the ratio $\Delta_{m}/\tilde{\Delta}_{c}$. As illustrated in Figs. 2(a)-2(c), each steady-state phase diagram exists three different regions corresponding to different phases, i.e., PSP, PSBP, and BP. It is worth noting that the closed system (i.e., excluding system dissipations) without drive (for example, standard Dicke model 2 ) can experience an equilibrium QPT, whose occurrence is often accompanied by a mutation of the ground state of the system. Different from the equilibrium QPT, the nonequilibrium QPT occurs in the driven-dissipative system, heralding the alteration of the steady state (rather than the ground state) of the system. In this work, the cavity-magnon system is driven-dissipative, and each point in the phase diagrams corresponds to stable states of the system [cf. Figs. 2(a)-2(c)]. In the PSP (PSBP), the system is in the steady state without (with) macroscopic magnon excitation. However, the system has two steady states in the BP, one without macroscopic magnon excitation and the other with macroscopic magnon excitation (cf. Fig. 3 and related discussions). In Fig. 2, the vertical black solid line ($G=G_{\text{c1}}$) is the boundary between PSP and BP in the case of $\Delta_{m}/\tilde{\Delta}_{a}\leq\beta|_{G=G_{\text{c1}}}$, and then the blue and red solid curves ($G=G_{\text{c2}}$) are the boundaries between PSP and PSBP for $\Delta_{m}/\tilde{\Delta}_{a}\geq\beta|_{G=G_{\text{c1}}}$, BP and PSBP for $\Delta_{m}/\tilde{\Delta}_{a}\leq\beta|_{G=G_{\text{c1}}}$, respectively. These manifest that the boundaries between different phases are determined by the critical driving strengths $G_{\text{c1}}$ and $G_{\text{c2}}$. In particular, the intersection point of three boundaries emerges when $G_{\text{c1}}=G_{\text{c2}}$ at $\Delta_{m}/\tilde{\Delta}_{a}=\beta|_{G=G_{\text{c1}}}$. More interestingly, we find that different spinning directions (the left or right) give rise to QPT at different critical driving strength. Compared with the stationary case (i.e., no spinning with $\Delta_{F}=0$), the areas of three phases have a remarkable change for $\Delta_{F}>0$ and $\Delta_{F}<0$. Therefore, this QPT are highly controllable and can be tuned by the driving direction (or the spinning direction) and the spinning speed of the resonator. As a result, the nonreciprocal QPT can be achieved in this system.
Figure 2: Steady-state phase diagram of the system varying with the reduced driving strength $G/\kappa$ and the detuning ratio $\Delta_{m}/\tilde{\Delta}_{c}$, in which PSP, PSBP, and BP represent the areas for the parity-symmetric phase, parity-symmetry-broken phase and bistable phase, respectively. The Fizeau shifts are $\Delta_{F}/\kappa=-0.3$ in (a), $\Delta_{F}/\kappa=0$ in (b), and $\Delta_{F}/\kappa=0.3$ in (c). The other parameters are $\Delta_{c}/\kappa=3$, $J/\kappa=2.5$, and $\gamma/\kappa=1.$
In Figs. 3(a)-3(c), we further display the dynamical behaviors of the scaled mean magnon number $\left|M\right|^{2}/\left(\gamma/K\right)$ to reveal the macroscopic magnon excitations for different phases. Here the parameter values for each phase have marked with different-colored dots in Fig. 2(b). When the system is in BP [see Fig. 3(a)], we find that both the solutions $\left|M\right|_{0}^{2}$ and $\left|M\right|_{+}^{2}$ are stable and the steady-state magnon occupation can be dominated by the initial condition of the system. When the system stabilizes to the steady state $\left|M\right|_{0}^{2}$ $(=0)$, the steady state has the even parity, i.e., the parity symmetry of the system is conserved. Conversely, since $\left|M\right|_{+}^{2}$ can be odd or even (corresponding to the odd or even parity), it is difficult to say that the parity of the steady state is odd or even in this case of $\left|M\right|_{+}^{2}$. For the steady state with $\left|M\right|_{+}^{2}$, the parity symmetry is broken. It is noticed that PSP and BP cannot be distinguished by only tracking the dynamics of the system with an initial state near zero since the solution $\left|M\right|_{0}^{2}$ is stable for both phases. In this case, the initial state should be set far away from zero to distinguish the two phases and observe the phase transition. From Fig. 3(b), we see that only the nontrivial solution $\left|M\right|_{+}^{2}$ is stable in PSBP, so that the macroscopic magnon excitation (i.e., $\left|M\right|_{+}^{2}$) can continuously varies and the parity symmetry of the system is broken in this situation. As described in Fig. 3(c), there is no macroscopic magnon excitation in PSP, that is to say, only the solution $\left|M\right|_{0}^{2}$ $(=0)$ is stable and then the system is parity-symmetric. Besides, an interesting feature is that the nontrivial solution $\left|M\right|_{-}^{2}$ is always unstable in the whole parameter range since the four eigenvalues of the matrix $\mathbf{U}$ in Eq. (12) have at least one nonnegative real part.
Figure 3: Dynamics of the normalized magnon number for three phases at the points marked by different-colored dots in Fig. 2(b), with (a) $G/\kappa=1.7$ and $\Delta_{m}/\Delta_{c}=1$, (b) $G/\kappa=2.3$ and $\Delta_{m}/\Delta_{c}=1.2$, (c) $G/\kappa=2.3$ and $\Delta_{m}/\Delta_{c}=0.4$. In (a)-(c), $\left\langle a\right\rangle_{t=0}/\sqrt{\gamma/K}=2+1.5i$ and $\left\langle b\right\rangle_{t=0}/\sqrt{\gamma/K}=2-1.5i$ for the solid curves, while $\left\langle a\right\rangle_{t=0}/\sqrt{\gamma/K}=0.05+0.05i$ and $\left\langle b\right\rangle_{t=0}/\sqrt{\gamma/K}=0.05-0.05i$ for the dashed curves. The other parameters are $J/\kappa=2.5$ and $\gamma/\kappa=1.$
[…]
Figure 4: Normalized steady-state magnon number $\left|M\right|^{2}/\left(\gamma/K\right)$ is plotted as the scaled driving strength $G/\kappa$ for three different Fizeau shifts $\Delta_{F}/\kappa=-0.3$, 00, and $0.3$. The solid curves correspond to the analytical results in Eq. (6) and the circles correspond to the numerical results obtained by Eq. (5). (a) $\Delta_{m}/\kappa=4$ and initial conditions $\left\langle a\right\rangle_{t=0}/\sqrt{\gamma/K}=\left\langle b\right\rangle_{t=0}/\sqrt{\gamma/K}=0.2+0.2i$, (b) $\Delta_{m}/\kappa=2.2$ and initial conditions $\left\langle a\right\rangle_{t=0}/\sqrt{\gamma/K}=\left\langle b\right\rangle_{t=0}/\sqrt{\gamma/K}=10+10i$. The other parameters are $\Delta_{c}/\kappa=3$, $J/\kappa=2.5$, and $\gamma/\kappa=1.$
Next, to quantitatively describe the nonreciprocal QPT, we introduce the isolation parameter 47a
=$\mathfrak{R}=\left\{\begin{array}[]{cc}0,\text{ }&\left|M\right|^{2}\left(\Delta_{F}>0\right){\small=}\left|M\right|^{2}\left(\Delta_{F}<0\right),\\ \left|\frac{\left|M\right|^{2}\left(\Delta_{F}<0\right)-\left|M\right|^{2}\left(\Delta_{F}>0\right)}{\left|M\right|^{2}\left(\Delta_{F}<0\right)+\left|M\right|^{2}\left(\Delta_{F}>0\right)}\right|,&\left|M\right|^{2}\left(\Delta_{F}>0\right){\small\neq}\left|M\right|^{2}\left(\Delta_{F}<0\right).\end{array}\right.$ (13)
For a QPT without the spinning resonator (i.e., a conventional reciprocal QPT), the isolation parameter is $\mathfrak{R}=0$. A nonzero $\mathfrak{R}$ denotes the emergence of nonreciprocity in the phase transition. The higher the isolation parameter $\mathfrak{R}$ is, the stronger the nonreciprocity of the QPT is. Especially, $\mathfrak{R}=1$ corresponds to an ideal nonreciprocal QPT. The isolation parameter $\mathfrak{R}$ as the function of the Fizeau shift $\left|\Delta_{F}\right|/\kappa$ is shown in Figs. 5(a) and 5(b) for the second-order and first-order phase transitions, respectively. As expected, the isolation parameter is $\mathfrak{R}=0$ for the QPT with a stationary resonator (i.e., $\Delta_{F}=0$). While it is clear that, not only the nonreciprocal QPT occurs in a remarkably broad parameter range, but also the region of the ideal nonreciprocal QPT enlarges with the increasing of the Fizeau shift $\left|\Delta_{F}\right|$. Further, the ideal nonreciprocal QPT exists obviously two boundaries, one boundary condition (the left black curve) satisfies $G=G_{\text{c2}}\left(\Delta_{F}<0\right)$ and the other one (the right black curve) satisfies $G=G_{\text{c2}}\left(\Delta_{F}>0\right)$ in Fig. 5(a). In the same way, there exists a similar characteristic in Fig. 5(b), where the boundary condition (the left black curve) obeys $G=G_{\text{c1}}\left(\Delta_{F}<0\right)$ and the other one (the right black curve) obeys $G=G_{\text{c1}}\left(\Delta_{F}>0\right)$. An interesting feature in Fig. 5 is that in the vicinity of the left black curses, the isolation parameter $\mathfrak{R}$ experiences a sudden change from zero to one. This is because both mean magnon numbers $\left|M\right|^{2}\left(\Delta_{F}<0\right)$ and $\left|M\right|^{2}\left(\Delta_{F}>0\right)$ are zero when the driving strength is less than the critical strength. However, when the driving strength is more than the critical strength, $\left|M\right|^{2}\left(\Delta_{F}<0\right)$ is nonzero and $\left|M\right|^{2}\left(\Delta_{F}>0\right)$ is still zero. In this situation, we see that the isolation parameter $\mathfrak{R}$ undergoes a gradually change from one to zero. It can be well understood that $\left|M\right|^{2}\left(\Delta_{F}<0\right)$ is nonzero and $\left|M\right|^{2}\left(\Delta_{F}>0\right)$ is zero when the driving strength is less than the critical strength. On the other hand, when the driving strength is more than the critical strength, both $\left|M\right|^{2}\left(\Delta_{F}<0\right)$ and $\left|M\right|^{2}\left(\Delta_{F}>0\right)$ are not only nonzero, but also gradually increasing with the driving strength increasing.
Figure 5: Isolation parameter $\mathfrak{R}$ versus the scaled Fizeau shift $\left|\Delta_{F}\right|/\kappa$ and normalized driving strength $G/\kappa$. The left and right black curves correspond to the boundary conditions in Eqs. (8) and (9), respectively. The parameters are the same as those in Fig. 2.
Finally, we reveal more characteristics of the nonreciprocal QPT through numerically simulating the behavior of the correlated fluctuations of the magnon operators $\delta m^{\dagger}$ and $\delta m$. As shown in Fig. 6, we plot the expectation value $\langle\delta m^{\dagger}\delta m\rangle$ of correlated fluctuation as a function of the scaled coupling strength $G/\kappa$ for different values of $\Delta_{F}$. There is a common characteristic in both Figs. 6(a) and 6(b), that is, the mean correlated fluctuation $\langle\delta m^{\dagger}\delta m\rangle$ is close to zero when the driving strength $G$ is away from the critical value, but Fig. 6(a) (Fig. 6(b)) is of a asymptotic (catastrophic) divergent behavior around the critical point, which is connected with the second-order QPT from PSP to PSBP (the first-order QPT from PSP to BP). From Fig. 6, we notice especially that spinning the resonator decreases the critical driving strength for $\Delta_{F}<0$ or increases it for $\Delta_{F}>0$, compared with the stationary-resonator case ($\Delta_{F}=0$). Besides, we find from Figs. 4 and 6 that the critical positions where the QPT occurs in Fig. 6(a) (Fig. 6(b)) completely agree with those in Fig. 4(a) (Fig. 4(b)) as we expected. The results give further evidences for the nonreciprocal QPT.
Figure 6: Mean correlated fluctuation $\langle\delta m^{\dagger}\delta m\rangle$ as a function of the scaled driving strength $G/\kappa$ for three different Fizeau shifts $\Delta_{F}/\kappa=-0.3,0,$and $0.3$. The other parameters are the same with Fig. 4.
IV Discussions and conclusions
Before concluding, we briefly discuss the experimental feasibility of the present proposal. In cavity magnonics, the decay rate $\kappa$ of the microwave cavity, the damping rate $\gamma$ of the magnon mode, as well as the coupling strength $J$ between the photon and magnon modes are of the order $1$ MHz (e.g., $\kappa/2\pi=2.04$ MHz, $\gamma/2\pi=1.49$ MHz, and $J/2\pi=8.17$ MHz in Ref. 67 ). In addition, the nonlinear coefficient $K$ of the magnon Kerr effect is of the order $1$ nHz 03a , and the thermodynamic limit (i.e., the weak nonlinearity limit) $\gamma/K\rightarrow\infty$ can be easily satisfied. By controlling the strength of the flux drive on a Josephson parametric amplifier, the amplitude $G$ of the parametric driven can be adjusted from 00 to $6$ MHz 68 . As for the nonreciprocity induced by the Fizeau shift, it is proportional to the spinning speed of the resonator, for instance, a spinning speed of $12.9$ kHz can bring about the Fizeau shift $\Delta_{F}$ $\approx 1.14$ MHz 69 . Therefore, our proposal is likely to be achieved in in cavity-magnon systems with current technologies.
In summary, we have theoretically studied the nonreciprocal QPT in the cavity magnonic system consisting of a YIG sphere placed in a spinning microwave resonator. By rotating the microwave resonator, we show that the introduced Sagnac effect can significantly modify the critical driving strengths for both second- and first-order QPTs and lead to a largely adjustable range of the critical driving strength. We further find that the critical driving strength of the phase transition relies on both the driving direction (or the rotation direction) and the spinning speed. In consequence, a highly tunable nonreciprocal QPT can be achieved based on the Sagnac effect. Our proposal provides an alternative route to achieve nonreciprocal quantum phase in microwave magnonical system and may find promising applications in designing nonreciprocal magnonic devices.
Acknowledgments
This work is supported by the National Science Foundation for Distinguished Young Scholars of the Higher Education Institutions of Anhui Province (Grant No. 2022AH020097). G. Q. Zhang is supported by the National Natural Science Foundation of China (Grant No. 12205069).
Note added. -In preparing our manuscript, we became aware of a similar work on nonreciprocal superradiant phase transitions published on Physical Review Letters 61 .